# Highly accurate simulations of asymmetric black-hole scattering and cross validation of effective-one-body models

### v1

@@ Abstract
The study of unbound binary–black-hole encounters provides a gauge-invariant approach to exploring strong-field gravitational interactions in two-body systems, which can subsequently inform waveform models for bound orbits. In this work, we present 60 new highly accurate numerical relativity (NR) simulations of black-hole scattering, generated using the Spectral Einstein Code (SpEC). Our simulations include 14 spin-aligned configurations, as well as 16 configurations with unequal masses, up to a mass ratio of 10. We perform the first direct comparison of scattering angles computed using different NR codes, finding good agreement. We compare our NR scattering angle results to the post-Minkowskian (PM)-based effective-one-body (EOB) closed-form models SEOB-PM and $w_{\rm EOB}$, finding less than 5% deviation except near the scatter-capture separatrix. Comparisons with the post-Newtonian-based EOB evolution models SEOBNRv5 and TEOBResumS-Dalí reveal that the former agrees within 8% accuracy with non-spinning NR results across most parameter ranges, whereas the latter matches similarly at lower energies but diverges significantly at higher energies. Both evolution EOB models exhibit increased deviations for spinning systems, predicting a notably different location of the capture separatrix compared to NR. Our key finding is the first measurement of disparate scattering angles from NR simulations due to asymmetric gravitational-wave emission. We compare these results to SEOB-PM models constructed to calculate the scattering angle of a single black hole in asymmetric systems.
@@ I Introduction
In order to obtain the speed and accuracy required for data analysis, we need to develop fast and accurate waveform models Abac et al. (2025); Afshordi et al. (2023); LIGO Scientific Collaboration and Virgo Collaboration and KAGRA Collaboration (2024). Effective-one-body (EOB) Buonanno and Damour (1999, 2000); Damour et al. (2000); Damour (2001); Buonanno et al. (2006, 2007) is one of the most successful approaches to modelling the dynamics and gravitational waveforms of binary systems. The EOB formalism is based on the idea of mapping the dynamics of a binary system to that of a single effective particle moving in a curved spacetime. This is obtained by deforming a Schwarzschild (or Kerr) Hamiltonian with terms from perturbative expansions (post-Newtonian Ramos-Buades et al. (2023); Gamba et al. (2022), post-Minkowskian Buonanno et al. (2024a); Damour et al. (2025), self-force van de Meent et al. (2023); Leather et al. (2025), etc.), and via calibration to NR data Pompili et al. (2023); Nagar et al. (2018). The resummed EOB Hamiltonian is evolved via Hamilton’s equations, with additional terms to account for the radiation reaction. Current state-of-the-art “evolution” [^1] Here, by evolution models we mean models obtained by solving Hamilton’s equations in time domain. EOB models fall into two families: SEOBNR Pompili et al. (2023); Ramos-Buades et al. (2023); van de Meent et al. (2023); Khalil et al. (2023); Gamboa et al. (2024); Buonanno et al. (2024a); Leather et al. (2025) and TEOBResumS Nagar et al. (2018, 2020a, 2020b); Gamba et al. (2022); Nagar et al. (2023); Albanesi et al. (2025a).
@@
As detector sensitivity improves, so must the accuracy of our waveform models to avoid systematic biases in parameter recovery due to modelling error, which would diminish the science return from the improved sensitivity Dhani et al. (2024); Pürrer and Haster (2020). Additionally, as detectors become sensitive to a broader range of physical properties (larger mass asymmetry, eccentricity, spin configurations), waveform models need to be updated to cover the relevant parameter space. Both the recent science case for the Einstein Telescope Abac et al. (2025) and the White Paper commissioned by the LISA Consortium Afshordi et al. (2023) call for accelerated progress on BBH waveforms and highlight the need for fresh ideas.
@@
One promising approach is based on the study of BH scattering encounters. The idea here is to use information of the high-energy unbound interactions to inform models of bound systems. An advantage of scattering encounters is that we can define the initial conditions and certain observables in terms of asymptotic quantities defined at past/future infinity. This allows us to avoid the local gauge ambiguities that plague comparisons of bound systems, where the initial conditions are defined at a finite separation. The scattering angle is a particularly useful asymptotic observable as it is directly related to the deflection of the BHs due to their gravitational interactions. As illustrated in Figure 1, variations in the initial conditions can significantly alter the black hole trajectories and, in turn, the resulting scattering angle.
@@
Recent state-of-the art PM results include the complete 5PM(1SF) scattering angle Driesse et al. (2024a, b), produced using the Worldline Quantum Field Theory formalism Mogull et al. (2021); Jakobsen et al. (2022). Here first self-force order (1SF) indicates that this result is valid only up to terms linear in the symmetric mass ratio $\nu$, defined in Eq. (1c). In this paper we also make use of 4PM non-spinning Dlapa et al. (2023a, b); Damgaard et al. (2023) and 5PM spinning results Jakobsen and Mogull (2022, 2023); Febres Cordero et al. (2023); Jakobsen et al. (2023a, b), which have been obtained using both worldlines and scattering amplitudes. Our PM counting convention used here is a physical one, and includes factors of $G$ arising from higher powers of spin. Other recent work includes a 7PM scattering angle at quartic order in spin on one BH Akpinar et al. (2025), and an extraction of the local-in-time part of the scattering angle relevant also for describing bound orbits Bini and Damour (2024); Dlapa et al. (2024, 2025).
While highly successful, the PM expansion becomes inaccurate as the separation between the bodies decreases and strong-field effects from the large curvature can no longer be neglected. Resummation techniques have incorporated information about the divergence of the scatter-capture separatrix to provide more faithful models across all separations Damour and Rettegno (2023); Long et al. (2024). EOB resummation shines in the context of a PM expansion because, as we shall review in Section III, an EOB model may be matched order-by-order to gauge-invariant scattering observables. The resulting “closed-form” EOB models can be used to make direct predictions for the scattering angle in the non-perturbative regime, without the need to fully evolve the system through Hamilton’s equations. Instead, the angle can be extracted by numerically evaluating a suitable integral.
The structure of the paper is as follows. In Sec. II we summarise the Spectral Einstein Code (SpEC) and the extensions we have made to the code to enable scattering encounters, detail our procedure to extract the scattering angle, and perform a direct comparison between scattering angle results from different NR codes. Section III reviews the EOB formalism and introduces the models used to calculate the scattering angle, namely the evolution model SEOBNRv5, and the closed-form models SEOB-PM and $w_{\rm EOB}$. We compare the scattering angle results from SpEC and the EOB models in Sec. IV for a variety of initial conditions including aligned spins, and unequal masses. We conclude in Sec. V with a summary of our findings and provide an outlook.
@@
Throughout this work, we use natural geometrised units with $G=c=1$. Our setup consists of two BHs with masses $m_{1}$ and $m_{2}$ with $m_{1}\geq m_{2}$. We denote the total mass, mass ratio and symmetric mass ratio by
@@
where $\mu$ represents the reduced mass. We utilise a rescaled version of the Arnowitt–Deser–Misner (ADM) energy, $E_{\rm ADM}$, defined by
@@
All BHs considered in our work have their spin vector parallel to the orbital angular-momentum vector, and are characterised by the projection of the spin onto the angular momentum direction, denoted by $S_{i},i=1,2$. We obtain the orbital angular momentum $L$ from the ADM angular momentum
@@
We define the dimensionless spins as $\chi_{i}=S_{i}/m_{i}^{2}$ which have a range $-1\leq\chi_{i}\leq 1$. For spins aligned with the orbital angular momentum, it is useful to define the following combinations
@@ II Numerical Relativity simulations
To enable accurate and efficient simulations of hyperbolic encounters, we extend and improve SpEC’s initial data routines, as well as evolution algorithm, as described in Sects. II.1 and II.2. We then perform a variety of NR simulations, as summarized in Sec. II.3. Section II.4 describes our techniques of extracting the scattering angle with error estimates, which we compare to results in the literature in Sec. II.5.
@@ II.1 Improvements: Initial data
As stated under 2. above, the initial data code requires $\Omega_{0}$ and $\dot{a}_{0}$ as part of its input parameters. However, for our scientific goals, it is much more convenient to be able to specify ADM energy and ADM angular momentum of the initial data sets. We included new root-finding functionality that adjusts $\Omega_{0}$ and $\dot{a}_{0}$ such that the final initial-data set achieves user-specified $E_{\rm ADM}$ and $J_{\rm ADM}$.
@@ II.2.1 Changes to the adaptive mesh refinement
Figure 2: Spacetime diagram of a $q=3$ SpEC scattering simulation. The solid blue and orange lines show the trajectories of the large and small BHs, respectively. The dashed lines represent the junk radiation emanating from the BH of the matching colour, and the black discs show when the junk from one BH reaches the other. In the red-shaded region, encompassing the propagation of junk radiation, AMR is inactive and the red dotted line shows when AMR is turned on. The line $t_{\rm branch}$ marks the start of the lower resolution runs.
@@ II.3 NR simulations
The parameter space of scattering simulations can be unambiguously defined in terms of quantities defined at spatial infinity $\Gamma$, $J_{\rm ADM}/(m_{1}m_{2})$, as well as intrinsic BH properties ($q$, $\chi_{i}$). The overall scale invariance of vacuum GR is represented by suitable rescalings to make all these parameters dimensionless. For the unequal-mass case, it is more convenient to rescale the asymptotic quantities to those that tend to the geodesic limit as $\nu\rightarrow 0$, such as $\gamma$ and $\ell$ defined in Eqs. (3) and (5) respectively. The SpEC simulations start with the BHs in the centre-of-mass frame at Cartesian coordinates $\vec{c}_{A}\approx(D_{0}/(1+q),0,0)$ and $\vec{c}_{B}\approx(-D_{0}\,q/(1+q),0,0)$ (GR corrections cause small deviations relative to these Newtonian values, see Ref. Ossokine et al. (2015)). We end the simulations when the BHs reach a separation of $D=1.4D_{0}$ on the outgoing leg. For all the SpEC simulations presented here, we use $D_{0}=250M$.
@@
Figure 3: Parameter space coverage of simulations: ‘SpEC $q\!=\!1,\chi_{i}\!=\!0$’ indicates equal-mass, non-spinning simulations presented here, and the large open circles denote sequences of SpEC simulations of unequal masses (up to $q\!=\!10$), or with non-zero spins (up to $\chi_{i}\!=\!0.60$). The grey solid lines show contours of constant scattering angle generated using SEOB-4PM, delineated by a fit of the scatter-capture separatrix (red dashed line) Kankani and McWilliams (2024). For context, the plot also shows parameters of equal-mass, non-spinning BBH scattering simulations performed with ETK Damour et al. (2014); Hopper et al. (2023); Rettegno et al. (2023); Swain et al. (2025) and GRA++ Albanesi et al. (2025b). Ref. Swain et al. (2025) also presents simulations extending up to $\gamma=1.96$, which are not shown on this plot.
@@ II.4 Extracting the scattering angle
The motion of each BH is tracked as the Cartesian coordinates of the centre of the excised regions in the initial centre-of-mass frame of the system. Each trajectory is split into incoming and outgoing legs and converted into polar $(r,\varphi)$ coordinates in the $xy$-plane.
@@
$\varphi=\varphi^{\infty}+\frac{\varphi^{(1)}}{D}+\frac{\varphi^{(2)}}{D^{2}}+\frac{\varphi^{(3)}}{D^{3}}+\ldots$ (9)
where $D$ is the separation of the BHs Damour et al. (2014); Damour and Rettegno (2023); Hopper et al. (2023); Rettegno et al. (2023); Albanesi et al. (2025b); Swain et al. (2025); Fontbuté et al. (2025a).
Alternatively, we can utilise the Keplerian relation between the separation and azimuthal angle
@@
where we use $+$ for the ingoing leg and -- for the outgoing leg.
@@
Figure 4: Calculation of the asymptotic scattering angle. The left panel shows the extrapolated outgoing angle obtained by fits over varying intervals $[d_{\rm min},350M]$. The Keplerian fit Eq. (11) yields more stable results than the polynomial fit Eq. (9). The right panel indicates the asymptotic scattering angle calculated from the fits in the left panel, and our choice of error bar. Both panels show three different resolutions (Lev1 through Lev3), which all give consistent results.
@@ II.5 Comparison between NR codes
Figure 5: Comparison of scattering angle extracted from SpEC and ETK Swain et al. (2025) simulations. All simulations start as equal mass, non-spinning binaries with $\Gamma=1.02264$. The top panel shows the relative scattering angle as a function of the initial angular momentum. The bottom panel shows the relative difference between the two codes. The SpEC values lie on $\delta\theta=0$ and have errors too small to be resolved on this scale, but are resolved in the inset, which shows a zoom-in of the six largest $\ell$ values. The blue shaded region on the lower plot demonstrates the expected difference in the angle due to a ±$\pm 10^{-5}$ difference in the initial energy. The vertical dotted lines show the first confirmed capture from each code.
@@
where $\theta_{1/2}$ are the scattering angles of each BH, as defined in Eq. (8). Both codes exhibit the same divergent behaviour as the BHs approach the scatter-capture separatrix. We note that the leftmost point ($\ell=4.3536$) only has a data point for ETK as the SpEC simulation produced a dynamical capture. In Table II of Ref. Swain et al. (2025) this run is marked as a potential dynamical capture due to the uncertainty in the scatter-capture separatrix and the fact that the BHs may merge if allowed to evolve to larger radii. Indeed, after the initial encounter, the SpEC simulation results in the BHs on a bound orbit with large apastron of $D=227.6M$ before merging during the secondary interaction. The ETK simulations are terminated at an outgoing separation of $D=180M$ and hence the capture would not have been observed.
@@
The values from both codes show good agreement with each other, differing by less than $1\%$ relative difference. However, the scattering angles reported by either code do not always lie within each other’s error bars. Both codes report the initial energy and angular momentum with an accuracy of approximately one part in $10^{5}$ (see Appendix A for a discussion of the SpEC accuracy). To demonstrate that this level of uncertainty is not responsible for the observed difference, we show the expected variation in the scattering angle resulting from a $10^{-5}$ change in the initial energy as the blue shaded region in the lower plot. This is obtained by varying the initial energy of the SEOB-4PM model (to be introduced in Sec. III.2) by ±$\pm 10^{-5}$ and computing the corresponding change in the scattering angle. The observed difference $\theta_{\textsc{ETK}}-\theta_{\textsc{SpEC}}$ at small $\ell$ is significantly larger than what can be attributed to variations in the initial conditions.

### v2

@@ Abstract
The study of unbound binary–black-hole encounters provides a gauge-invariant approach to exploring strong-field gravitational interactions in two-body systems, which can subsequently inform waveform models for bound orbits. In this work, we present 60 new highly accurate numerical relativity (NR) simulations of black-hole scattering, generated using the Spectral Einstein Code (SpEC). Our simulations include 14 spin-aligned configurations, as well as 16 configurations with unequal masses, up to a mass ratio of 10. We perform the first direct comparison of scattering angles computed using different NR codes, finding good agreement. We compare our NR scattering angle results to the post-Minkowskian (PM)-based effective-one-body (EOB) closed-form models SEOB-PM and $w_{\rm EOB}$, finding less than 5% deviation except near the scatter-capture separatrix. Comparisons with the post-Newtonian-based EOB evolution models SEOBNRv5 and TEOBResumS-Dalí reveal that the former agrees within 8% accuracy with non-spinning NR results across most parameter ranges, whereas the latter matches similarly at lower energies but diverges significantly at higher energies. Both evolution EOB models exhibit increased deviations for spinning systems, predicting a notably different location of the capture separatrix compared to NR. Our key result is the first measurement of disparate scattering angles from NR simulations due to asymmetric gravitational-wave emission. We compare these results to SEOB-PM models constructed to calculate the scattering angle of a single black hole in asymmetric systems.
@@ I Introduction
In order to obtain the speed and accuracy required for data analysis, we need to develop fast and accurate waveform models Abac et al. (2025); Afshordi et al. (2023); LIGO Scientific Collaboration and Virgo Collaboration and KAGRA Collaboration (2024). Effective-one-body (EOB) Buonanno and Damour (1999, 2000); Damour et al. (2000); Damour (2001); Buonanno et al. (2006, 2007) is one of the most successful approaches to modeling the dynamics and gravitational waveforms of binary systems. The EOB formalism is based on the idea of mapping the dynamics of a binary system to that of a single effective particle moving in a curved spacetime. This is obtained by deforming a Schwarzschild (or Kerr) Hamiltonian with terms from perturbative expansions (post-Newtonian Ramos-Buades et al. (2023); Gamba et al. (2022), post-Minkowskian Buonanno et al. (2024a); Damour et al. (2025), self-force van de Meent et al. (2023); Leather et al. (2025), etc.), and via calibration to NR data Pompili et al. (2023); Nagar et al. (2018). The resummed EOB Hamiltonian is evolved via Hamilton’s equations, with additional terms to account for the radiation reaction. Current state-of-the-art “evolution” [^1] Here, by evolution models we mean models obtained by solving Hamilton’s equations in time domain. EOB models fall into two families: SEOBNR Pompili et al. (2023); Ramos-Buades et al. (2023); van de Meent et al. (2023); Khalil et al. (2023); Gamboa et al. (2024); Buonanno et al. (2024a); Leather et al. (2025) and TEOBResumS Nagar et al. (2018, 2020a, 2020b); Gamba et al. (2022); Nagar et al. (2023); Albanesi et al. (2025a).
@@
As detector sensitivity improves, so must the accuracy of our waveform models to avoid systematic biases in parameter recovery due to modeling error, which would diminish the science return from the improved sensitivity Dhani et al. (2025); Pürrer and Haster (2020). Additionally, as detectors become sensitive to a broader range of physical properties (larger mass asymmetry, eccentricity, spin configurations), waveform models need to be updated to cover the relevant parameter space. Both the recent science case for the Einstein Telescope Abac et al. (2025) and the White Paper commissioned by the LISA Consortium Afshordi et al. (2023) call for accelerated progress on BBH waveforms and highlight the need for fresh ideas.
@@
One promising approach is based on the study of BH scattering encounters. The idea here is to use information of the high-energy unbound interactions to inform models of bound systems Damour (2016, 2018). An advantage of scattering encounters is that we can define the initial conditions and certain observables in terms of asymptotic quantities defined at past/future infinity. This allows us to avoid the local gauge ambiguities that plague comparisons of bound systems, where the initial conditions are defined at a finite separation. The scattering angle is a particularly useful asymptotic observable as it is directly related to the deflection of the BHs due to their gravitational interactions. As illustrated in Figure 1, variations in the initial conditions can significantly alter the black hole trajectories and, in turn, the resulting scattering angle.
@@
Recent state-of-the art PM results include the complete 5PM(1SF) scattering angle Driesse et al. (2024, 2025), produced using the Worldline Quantum Field Theory formalism Mogull et al. (2021); Jakobsen et al. (2022). Here first self-force order (1SF) indicates that this result is valid only up to terms linear in the symmetric mass ratio $\nu$, defined in Eq. (1c). In this paper we also make use of 4PM non-spinning Dlapa et al. (2023a, b); Damgaard et al. (2023) and 5PM spinning results Jakobsen and Mogull (2022, 2023); Febres Cordero et al. (2023); Jakobsen et al. (2023a, b), which have been obtained using both worldlines and scattering amplitudes. Our PM counting convention used here is a physical one, and includes factors of $G$ arising from higher powers of spin. Other recent work includes a 7PM scattering angle at quartic order in spin on one BH Akpinar et al. (2025), and an extraction of the local-in-time part of the scattering angle relevant also for describing bound orbits Bini and Damour (2024); Dlapa et al. (2024, 2025).
While highly successful, the PM expansion becomes inaccurate as the separation between the bodies decreases and strong-field effects from the large curvature can no longer be neglected. Resummation techniques have incorporated information about the divergence of the scatter-capture separatrix to provide more faithful models across all separations Damour and Rettegno (2023); Long et al. (2024). EOB resummation shines in the context of a PM expansion because, as we shall review in Section III, an EOB model may be matched order-by-order to gauge-invariant scattering observables Damour (2016, 2018). The resulting “closed-form” EOB models can be used to make direct predictions for the scattering angle in the non-perturbative regime, without the need to fully evolve the system through Hamilton’s equations. Instead, the angle can be extracted by numerically evaluating a suitable integral.
The structure of the paper is as follows. In Sec. II we summarize the Spectral Einstein Code (SpEC) and the extensions we have made to the code to enable scattering encounters, detail our procedure to extract the scattering angle, and perform a direct comparison between scattering angle results from different NR codes. Section III reviews the EOB formalism and introduces the models used to calculate the scattering angle, namely the evolution model SEOBNRv5, and the closed-form models SEOB-PM and $w_{\rm EOB}$. We compare the scattering angle results from SpEC and the EOB models in Sec. IV for a variety of initial conditions including aligned spins, and unequal masses. We conclude in Sec. V with a summary of our findings and provide an outlook.
@@
Throughout this work, we use natural geometrised units with $G=c=1$. Our setup consists of two BHs with masses $m_{1}$ and $m_{2}$ with $m_{1}\geq m_{2}$. We denote the total mass, mass ratio, symmetric and asymmetric mass ratios by
@@
$\delta$ $=\frac{m_{1}-m_{2}}{M},$ (1d)
where $\mu$ represents the reduced mass. We utilize a rescaled version of the Arnowitt–Deser–Misner (ADM) energy, $E_{\rm ADM}$, defined by
@@
All BHs considered in our work have their spin vector parallel to the orbital angular-momentum vector, and are characterized by the projection of the spin onto the angular momentum direction, denoted by $S_{i},i=1,2$. We obtain the orbital angular momentum $L$ from the ADM angular momentum
@@
We define the dimensionless spins as $\chi_{i}=a_{i}/m_{i}=S_{i}/m_{i}^{2}$ which have a range $1\leq\chi_{i}\leq 1$. For spins aligned with the orbital angular momentum, it is useful to define the following combinations
@@ II Numerical Relativity simulations
To enable accurate and efficient simulations of hyperbolic encounters, we extend and improve SpEC’s initial data routines, as well as evolution algorithm, as described in Secs. II.1 and II.2. We then perform a variety of NR simulations, as summarized in Sec. II.3. Section II.4 describes our techniques of extracting the scattering angle with error estimates, which we compare to results in the literature in Sec. II.5.
@@ II.1 Improvements: Initial data
As stated under 2. above, the initial data code requires $\Omega_{0}$ and $\dot{a}_{0}$ as part of its input parameters. However, for our scientific goals, it is much more convenient to be able to specify ADM energy and ADM angular momentum of the initial data sets. We included new root-finding functionality that adjusts $\Omega_{0}$ and $\dot{a}_{0}$ such that the final initial-data set achieves user-specified $E_{\rm ADM}$ and $J_{\rm ADM}$. Details of this routine are presented in Secs. II D and E of Ref. Mendes et al. (2025).
@@ II.2.1 Changes to the adaptive mesh refinement
Figure 2: Spacetime diagram of a $q=3$ SpEC scattering simulation. The solid blue and orange lines show the trajectories of the large and small BHs, respectively. The dashed lines represent the junk radiation emanating from the BH of the matching color, and the black discs show when the junk from one BH reaches the other. In the red-shaded region, encompassing the propagation of junk radiation, AMR is inactive and the red dotted line shows when AMR is turned on. The line $t_{\rm branch}$ marks the start of the lower resolution runs. We show the value of the outer boundary of the computational domain, $R_{\rm bdry}$ for reference.
@@ II.3 NR simulations
The parameter space of scattering simulations can be unambiguously defined in terms of quantities defined at spatial infinity $\Gamma$, $J_{\rm ADM}/(m_{1}m_{2})$, as well as intrinsic BH properties ($q$, $\chi_{i}$). The overall scale invariance of vacuum GR is represented by suitable rescalings to make all these parameters dimensionless. For the unequal-mass case, it is more convenient to rescale the asymptotic quantities to those that tend to the geodesic limit as $\nu\rightarrow 0$, such as $\gamma$ and $\ell$ defined in Eqs. (3) and (5) respectively. The SpEC simulations start with the BHs in the center-of-mass frame at Cartesian coordinates $\vec{c}_{A}\approx(D_{0}/(1+q),0,0)$ and $\vec{c}_{B}\approx(-D_{0}\,q/(1+q),0,0)$ (GR corrections cause small deviations relative to these Newtonian values, see Ref. Ossokine et al. (2015)). We end the simulations when the BHs reach a separation of $D=1.4D_{0}$ on the outgoing leg. For all the SpEC simulations presented here, we use $D_{0}=250M$.
@@
Figure 3: Parameter space coverage of simulations: ‘SpEC $q\!=\!1,\chi_{i}\!=\!0$’ indicates equal-mass, non-spinning simulations presented here, and the large open circles denote sequences of SpEC simulations of unequal masses (up to q=10q\!=\!10), or with non-zero spins (up to $\chi_{i}\!=\!0.60$). The grey solid lines show contours of constant scattering angle generated using SEOB-4PM, delineated by a fit of the scatter-capture separatrix (red dashed line) Kankani and McWilliams (2024). For context, the plot also shows parameters of equal-mass, non-spinning BBH scattering simulations performed with ETK Damour et al. (2014); Hopper et al. (2023); Rettegno et al. (2023); Swain et al. (2025) and GRA++ Albanesi et al. (2025b). Ref. Swain et al. (2025) also presents simulations extending up to $\gamma=1.96$, which are not shown on this plot.
@@ II.4 Extracting the scattering angle
The motion of each BH is tracked as the Cartesian coordinates of the center of the excised regions in the initial center-of-mass frame of the system. Each trajectory is split into incoming and outgoing legs and converted into polar $(r,\varphi)$ coordinates in the $xy$-plane.
@@
$\varphi=\varphi^{\infty}+\frac{\varphi^{(1)}}{D}+\frac{\varphi^{(2)}}{D^{2}}+\frac{\varphi^{(3)}}{D^{3}},$ (9)
where $D$ is the separation of the BHs. Previous works Damour et al. (2014); Damour and Rettegno (2023); Hopper et al. (2023); Rettegno et al. (2023); Albanesi et al. (2025b); Swain et al. (2025); Fontbuté et al. (2025a) varied the order of the polynomial whereas here we use a cubic polynomial in $1/D$, which we found to provide a good fit to the data without overfitting.
Alternatively, we can utilize the Keplerian relation between the separation and azimuthal angle
@@
where we use ++ for the ingoing leg and - for the outgoing leg.
@@
Figure 4: Calculation of the asymptotic scattering angle. The left panel shows the extrapolated outgoing angle obtained by fits over varying intervals $d_{\rm out},350M]$. The Keplerian fit Eq. (11) yields more stable results than the polynomial fit Eq. (9). The right panel indicates the asymptotic scattering angle calculated from both fits in the left panel, and our choice of error bar. Both panels show three different resolutions (Lev1 through Lev3), which all give consistent results.
@@ II.5 Comparison between NR codes
Figure 5: Comparison of scattering angle extracted from SpEC and ETK Swain et al. (2025) simulations. All simulations start as equal mass, non-spinning binaries with $\Gamma=1.02264$. The top panel shows the relative scattering angle as a function of the initial angular momentum. The bottom panel shows the relative difference between the two codes. The SpEC values lie on $\delta\theta=0$ with the error bars being the errors of the upper panel rescaled by the value of the angle. These errors are too small to be resolved on this scale, but are resolved in the inset, which shows a zoom-in of the six largest $\ell$ values. The blue shaded region on the lower plot demonstrates the expected difference in the angle due to a $\pm 10^{-5}$ difference in the initial energy. The vertical dotted lines show the first confirmed capture from each code.
@@
where $\theta_{1/2}$ are the scattering angles of each BH, as defined in Eq. (8). Both codes exhibit the same divergent behavior as the BHs approach the scatter-capture separatrix. We note that the leftmost point ($\ell=4.3536$) only has a data point for ETK as the SpEC simulation produced a dynamical capture. In Table II of Ref. Swain et al. (2025) this run is marked as a potential dynamical capture due to the uncertainty in the scatter-capture separatrix and the fact that the BHs may merge if allowed to evolve to larger radii. Indeed, after the initial encounter, the SpEC simulation results in the BHs on a bound orbit with large apastron of $D=227.6M$ before merging during the secondary interaction. The ETK simulations are terminated at an outgoing separation of $D=180M$ and hence the capture would not have been observed.
@@
The values from both codes show good agreement with each other, differing by less than $1\%$ relative difference. However, the scattering angles reported by either code do not always lie within each other’s error bars. Both codes report the initial energy and angular momentum with an accuracy of approximately one part in $10^{5}$ (see Appendix A for a discussion of the SpEC accuracy). To demonstrate that this level of uncertainty is not responsible for the observed difference, we show the expected variation in the scattering angle resulting from a $10^{-5}$ change in the initial energy as the blue shaded region in the lower plot. This is obtained by varying the initial energy of the SEOB-4PM model (to be introduced in Sec. III.2) by $\pm 10^{-5}$ and computing the corresponding change in the scattering angle. The observed difference $\theta_{\textsc{ETK}}-\theta_{\textsc{SpEC}}$ at small $\ell$ is significantly larger than what can be attributed to variations in the initial conditions.